OpenStudy (anonymous):

@sasogeek please point out any errors you see please :) -x^4-x^2+9=0

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] a=-x^4 b=x^2 c=9

OpenStudy (anonymous):

the discriminant is b^2-4ac. So, the discriminant is 12^2-4(4)(9) Which is 144-144 And comes out to be 0.

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

first ima multiply the equation by -1 to make the variables positive then change the variable to u u^2=x^4 u=s^2

OpenStudy (anonymous):

so -1(-x^4-x^2+9) x^4+x^2-9=0 x^4=u^2 x^2=u so now the equation is \[u^2+u-9=0\]

OpenStudy (anonymous):

a=u^2 b=u c=-9

OpenStudy (anonymous):

ok my work here is done

OpenStudy (anonymous):

or 1,1,-9 forgot to substitute in the 1's

OpenStudy (anonymous):

\[x=\frac{ -1\pm \sqrt{1^2}-4(1)(-9) }{ 2(1) }\] \[x=\frac{ -1 \pm \sqrt{1+36} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ -1 \pm \sqrt{37} }{ 2 }\] This is all of this and btw the first equation in the last reply is suppose to be sqrt over all but -1 +- on top row

OpenStudy (sasogeek):

you're making sense for the most part but idk for sure if you're fully correct. i've lost touch with a lot of my math cos my academic career has taken a different path. i believe you're on the right path though. I'll ask a friend to confirm your work :)

OpenStudy (anonymous):

now i just solved for x and need x^2 also \[x=\frac{ -1+\sqrt{37} }{ 2 }\] and \[x=\frac{ -1+\sqrt{37} }{ 2 }\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

ima continue to work so i can get u^2 this is where i confuse myself at times

OpenStudy (anonymous):

so those problems are u which is x^2 and we need u^2 also or x^4 to do that im thinking raise everything to the 2nd power \[x=\frac{ -2\pm \sqrt{74} }{ 4 }\] so the last 2 solutions are \[(x=\frac{ -1 \pm \sqrt{37}}{ 2 })^2\]\[x^2=\frac{ 1 \pm 37 }{ 4 }\] \[x^2=\frac{ 38 }{ 4 }\] ???

OpenStudy (aum):

-x^4-x^2+9=0 multiply throughout by -1: x^4 + x^2 - 9 = 0 Let u = x^2. Then u^2 = x^4 u^2 + u - 9 = 0 a = 1, b = 1, c = -9 plug it into the quadratic formula and solve for u. Then take the square root to solve for x.

OpenStudy (anonymous):

i know all that im just practicing so i dont get as confused while solving

OpenStudy (anonymous):

and would like to know when i mess up

ganeshie8 (ganeshie8):

your work for solving \(\large u\) looks good and the answer is correct ! \[\large \begin{align}\\u =x^2 &= \dfrac{1}{2}(-1\pm\sqrt{37}) \\~\\ x&= \pm \sqrt{ \dfrac{1}{2}(-1\pm\sqrt{37})}\end{align}\] you may simplify further if possible or leave it here...

OpenStudy (anonymous):

can you explain how you got the 1/2 please? did you multiply from the denominator?

ganeshie8 (ganeshie8):

that expression is same as yours, see if this looks better : \[\large \begin{align}\\u =x^2 &= \dfrac{-1\pm\sqrt{37}}{2} \\~\\ x&= \pm \sqrt{ \dfrac{-1\pm\sqrt{37}}{2}}\end{align} \]

ganeshie8 (ganeshie8):

\[\large \dfrac{-1\pm\sqrt{37}}{2} \] is same as \[\large \dfrac{1}{2}(-1\pm\sqrt{37}) \]

OpenStudy (anonymous):

oh i see now thanks and that would be the final answer? no need to go any further?

ganeshie8 (ganeshie8):

I prefer the latter form thats ll... :)

ganeshie8 (ganeshie8):

Looks good to me !

ganeshie8 (ganeshie8):

unless if there is a way to simplify the radical : \[\large x= \pm \sqrt{ \dfrac{-1\pm\sqrt{37}}{2}}\] sometimes we can simplify radicals like these, but i don't think it is possible here... so you're done mostly !

OpenStudy (anonymous):

did i help

OpenStudy (anonymous):

ok thanks ganeshie :)

OpenStudy (aum):

There are 4 roots to the original equation: two real and two complex. The above is a compact way of expressing the 4 roots.